# bibliography.bib Relevant, hidden, and frustrated information in high-dimensional analyses of complex dynamical systems with internal noise

### v3

@@ Abstract
Extracting from trajectory data meaningful information to understand complex systems might be non-trivial. High-dimensional analyses are typically assumed to be desirable, if not required, to prevent losing important information. However, to what extent such high-dimensionality is really needed/beneficial often remains unclear. Here we challenge such a fundamental general problem. As first representative cases of a system with internal dynamical complexity, we study atomistic molecular dynamics trajectories of liquid water and ice coexisting in dynamical equilibrium in correspondence of the solid/liquid transition temperature. To attain an intrinsically high-dimensional analysis, we use as an example the Smooth Overlap of Atomic Positions (SOAP) descriptor, obtaining a large dataset containing 576-dimensional SOAP vectors that we analyze in various ways. Remarkably, our results demonstrate how the time-series data contained in one single SOAP dimension accounting only for ¡0.001 of the total variance (thus considered negligible and discarded in typical variance-based dimensionality reduction approaches) is as much, and even more descriptive than the entire high-dimensional dataset. Adding more dimensions to the analysis is found detrimental due to ”frustrated information” phenomena that are proven general and are observed also in the analysis of, e.g., experimental trajectories of collective convolutions of active microparticles, here used as a second case study, performed with a combination of physics-inspired descriptors. This work explicitly demonstrates how high-dimensional analyses are not necessarily better than low-dimensional ones. These results also underline the importance of considering the quality (relevance) over the quantity of information, especially when dealing with noisy high-dimensional datasets, where the noise of one dimension may have higher statistical weight than the relevant information contained into another.
@@ Introduction
A common assumption when studying complex systems, about which little is known a priori, is the need to perform high-dimensional analyses to avoid losing information due to prior assumptions or incomplete characterization. However, selecting a set of physically relevant descriptors capable of providing a complete characterization of a specific system requires a deep knowledge of the system itself, which is not always available. Moreover, managing high-dimensional analyses and extracting meaningful physical insights from them is often challenging. Fundamental questions arise: What is the intrinsic dimension of the dataset [facco2017estimating, dinoia2024beyond]? Which components are the most relevant?
@@
Similar to other descriptors, such as atomic cluster expansions (ACE) [drautz2019atomic] or N-body iterative contraction of equivariants (NICE) [nigam_recursive_2020], SOAP provides high-dimensional mathematical representations of local density and order/disorder. Recently, SOAP has been successfully applied to investigate complex molecular systems, including liquid-solid and liquid-liquid coexistence in aqueous systems [Donkor2024spatavg, offeidanso2022, Ansari2020, Martino2024], and non-trivial dynamics and local disorder events in metal surfaces and nanoparticles [Caruso2023tsoap, Caruso2024, crippa2023machine, Rapetti2023, Cioni2024, perrone2024], and ionic behaviors in dense environments [Lionello2022].
@@
This finding raises several important questions: To what extent is high-dimensionality truly necessary to capture the physics of these systems? How can a high-dimensional analysis be less informative than its dimensionally-reduced counterparts? What type of information remain hidden, where is it located, and how can it be extracted? Is the information contained in the orthogonal dimensions of a high-dimensional feature space always additive, does adding a second dimension always contribute new information? These questions are particularly central when using intrinsically high-dimensional descriptors such as SOAP or ACE [drautz2019atomic], but they also apply to any combination of descriptors, whether physics-inspired, abstract, or general. At their core, these questions challenge a fundamental issue in physics: how to effectively extract meaningful information from high-dimensional analyses, a problem that is central in many fields.
@@
Typically, high-dimensional dataset extracted from trajectories, are studied as temporally independent under an ergodic assumption, and the identified high density peaks detect dominant/recurrent SOAP domains [Capelli2022, Gardin2022]. However, we have recently shown that significant information may be hidden in the temporal correlations of the data – such as local, sparse but dominant fluctuations, important transitions, and more – which are overlooked in pattern recognition approaches which neglect time correlations. We also developed an efficient method to extract and classify such information, systematically distinguishing statistically relevant fluctuations from noise in any type of time-series data [Becchi2024layer].
@@
In this work, we tackle the longstanding and challenging problem of extracting meaningful information from multidimensional datasets through systematic analysis. As a prototypical example of an information-rich dataset, we utilize data extracted from atomistic molecular dynamics (MD) simulation of a periodic box where liquid water and ice coexist in dynamical equilibrium at the melting temperature [Caruso2023tsoap, crippa2023machine, Becchi2024layer, Caruso2024, Martino2024]. This system, which has previously been shown to exhibit nontrivial internal complexity, is explored here by computing 576-dimensional SOAP spectra for each of the 2048 water molecules in the system every ps (sampling interval) over 50 ns of MD trajectory. This results in a dataset of 576-dimensional SOAP spectra, which we analyze both by ignoring and considering the time correlations (i.e., pattern recognition vs. time-series analyses).
@@
We demonstrate that dimensionality reduction methods based on variance, such as Principal Component Analysis (PCA), often fail when dealing with noisy data. Remarkably, we show that time-series data contained in a single SOAP component, which accounts for less than of the total variance (and thus typically discarded in PCA), provides more information than the entire high-dimensional dataset. Moreover, our results reveal that adding more (noisy) components to this single one can be not only non-beneficial but also detrimental to the analysis, a phenomenon we describe as “noise-frustrated information”. Finally, we show that such effects are not exclusive to SOAP, but can arise in any noisy multidimensional analysis. The findings presented here challenge classical paradigms in data analysis, highlighting the critical importance of focusing on information quality (relevance) rather than quantity to prevent information loss and misinterpretation.
@@ Results
The extraction of information from such a simulated system involves three main steps: (i) acquisition of raw data, which consist in the sequence of the particles’ coordinates acquired every ps, over ns of MD simulation; (ii) selection of an appropriate descriptor to convert the trajectory of each molecule – given by the time-series data – into a dataset ; (iii) analysis of the dataset to extract relevant information on the system’s physics (see Figure 1b). The information that can be effectively collected from the system depends on the choice of the descriptor in step (ii). The correct choice of descriptor is crucial for two main reasons. First, different descriptors capture distinct aspects of the system’s behavior: for example, a structural descriptor would be useless to study purely dynamical features [crippa2023machine]. Second, each descriptor has an intrinsic signal-to-noise ratio, which influences the clarity and reliability of the extracted information [Martino2024]. When the physics of a system is unknown, selecting a general, agnostic – possibly intrinsically high-dimensional – descriptor would be desirable in order to minimize the risk of introducing assumptions that could bias the analysis or result in incomplete information.
@@
It is particularly noteworthy that, in Figure 2a, four PCs were required to reach of the cumulative variance, implying that the dataset was approximately four-dimensional. However, the local noise reduction allows PC1 alone to account for nearly the entire variance of the dataset, indicating that the “relevant” information, without noise, is essentially mono-dimensional. This suggests that the high dimensionality observed in the original dataset is primarily a reflection of noise, not of the inherent complexity of the system.
@@
Information quantity vs. information quality in noisy dataset
@@
Notably, all our evidences demonstrated that this phenomenon is general and can occur independently on the number of dimensions, their individual noise levels (native vs. denoised), or the dimensionality reduction approach employed in the analysis. To further illustrate this point, Supporting Figure SI7 shows the manifold representations of the SOAP and SOAP datasets (computed using UMAP [mcinnes2018umap, Donkor2024spatavg]). This approach operates directly on the full high-dimensional dataset – meaning combining all component information “correctly” – and captures comparable physically relevant information to the one detected by the single component (Figure 4f, g). This confirms that discriminating the components by their variance is not necessarily the best approach, and better metrics should be used to extract the relevant information [Ebrahimi1999].
Information frustration effects are also observed when additional data are added in time (by increasing the sampling frequencies) rather than in space (by adding SOAP components). To investigate the effects of different sampling frequencies, we compared the results obtained analyzing the MD trajectories (raw data) sampled every ps (Figures 2,3,4) with those obtained by increasing or decreasing the sampling frequency. We therefore resampled the 50 ns MD trajectory at intervals of 4 ps, 20 ps, 80 ps, and 100 ps. The expectation is that Onion Clustering, which extracts relevant information about physical events in the system (for instance, the detectability of the ice/water interface environment up to a certain ), should maintain or enhance its ability to resolve this environment when using finer sampling ( ps). Conversely, coarser sampling ( ps) might compromise the resolution, potentially preventing proper detection of the interface. However, the results of these analyses reveal issues analogous to those encountered when adding more spatial information. As shown in Figure 5b, the black line represents the reference case ( ps): the third environment – the ice/water interface – is detectable within the resolution range 2 ns ns in the PC1 time-series, as seen in Figure 2c. Increasing the sampling time to 4 ps or 20 ps produces striking effects, as the temporal window within which the interface is detected shifts backward in the time-series. For a 20 ps sampling interval (which doubles the number of frames compared to the reference), the detectable temporal window for the interface shifts to 1 ns ns (green data). With a 4 ps sampling interval (a tenfold increase in numbers of analyzed frames), the interface is detectable only between 200 ps and 300 ps, a much shorter period compared to the nanosecond-scale range observed in other cases.
@@
Generality of the concept: analysis of experimental dynamical systems
@@
In the tests of Figures 1-5 we used SOAP as a representative example of a high-dimensional descriptor, providing high-dimensional datasets. However, the same phenomena can occur independently of the specific system or the set of descriptors/variables selected. To illustrate the generality of these concepts, we analyzed a completely different system, characterized by distinct scales and unique dynamics. In this case, we examined time-series obtained from experimentally-acquired trajectories [Liu2021quinckerollers, Becchi2024layer]. The system involves the two-dimensional collective movements of colloidal particles, known as Quincke rollers. Readers seeking more details on the system are referred to Ref. [Liu2021quinckerollers, Becchi2024layer, Caruso2024]. Briefly, Quincke rollers are dielectric polystyrene particles (m in diameter) suspended in a weakly conductive fluid and confined within a two-dimensional cell. Under the influence of a weak vertical DC electric field, these particles can exhibit complex collective motions, such as vortexes and waves. Here, we analyze the trajectory we resolved for one example Quincke rollers system [Becchi2024layer], consisting of 6921 particles moving within a m cell over a real-time duration of 0.25 s (310 frames in total). During this interval, a collective density wave is seen to propagate from left to right in the microscopy field (see Figure 6a-b).
@@
For generality purposes, this analysis uses two human-based descriptors. Recently, we have demonstrated that integrating static descriptors, such as SOAP, with dynamic ones, such as LENS can enrich high-dimensional analyses [crippa2023detecting]. In the same spirit – and reflecting common practice in studying collective dynamical events in active matter [peruani2008mean] – we selected two descriptors: (i) the minimum neighbor distance , as a proxy for the local particle density, and (ii) the local alignment of the particle velocity . This choice results in a bi-dimensional analysis (details are provided in the Methods section). We analyzed the denoised mono- and bi-dimensional time-series of the 6921 particles. Using Onion Clustering, was identified as a less informative descriptor (Figure 6c). From its time-series, two environments are detected up to ms: one representing the wave’s core (having higher density) and the other identifying static particles (with lower density). In contrast, the particles’ local velocity alignment (Figure 6d) is found to be more informative, capturing the system’s collective motion and identifying three distinct environments for resolutions higher than ms, and two environments for ms. When three clusters are detected, Onion Clustering distinguishes the wave’s core (Figure 6d in fuchsia) and two regions (cyan) where particles transition between static and dynamic states around the wave’s core. These regions are shown in the snapshot in Figure 6d, alongside with the static particles’ environment (dark violet). For ms, Onion Clustering fails to resolve any statistically distinct dynamical cluster, indicating that, in the system, all particles change environment at list once during the trajectory. The residence time of particles in any cluster is therefore less than 140 ms.
@@ Discussion
As a prototypical case study, we analyzed Smooth Overlap of Atomic Position (SOAP) spectra [bartok2013representing] extracted from MD simulations of coexisting ice and liquid water at dynamic equilibrium at the solid/liquid transition temperature. Typically, SOAP analyses of such systems involve dimensionality reduction methods, such as Principal Component Analysis (PCA), followed by clustering based on a combination of principal components (PCs) that account for high cumulative variance. However, approaches like this ignore the time-correlation in the data; this can result in significant information loss and limit the ability to capture key dynamic events [Becchi2024layer]. To maximize the information extraction, we opted to analyze the SOAP data in time-series using Onion Clustering [Becchi2024layer], which allows resolving the information within individual SOAP PC1 and PC2 time-series, as well as their combination in bi-variate time-series. Our results demonstrate that two-dimensional analysis systematically lead to information loss, independently of whether the time-series data are noisy or denoised [Donkor2024spatavg]. This phenomenon persists even when analyzing two of the original SOAP components, such as 63 and 237 (see Figures 4-5), and thus regardless of the chosen dimensionality reduction approach. We call this effect “frustrated information”, wherein relevant information captured in one dimension manifest as noise in another one.

### v5

@@ Abstract
Extracting from trajectory data meaningful information to understand complex molecular systems might be non-trivial. High-dimensional analyses are typically assumed to be desirable, if not required, to prevent losing important information. But to what extent such high-dimensionality is really needed/beneficial often remains unclear. Here we challenge such a fundamental general problem. As a representative case of a system with internal dynamical complexity, we study atomistic molecular dynamics trajectories of liquid water and ice coexisting in dynamical equilibrium at the solid/liquid transition temperature. To attain an intrinsically high-dimensional analysis, we use as an example the Smooth Overlap of Atomic Positions (SOAP) descriptor, obtaining a large dataset containing 576-dimensional SOAP vectors that we analyze in various ways. Our results demonstrate how the time-series data contained in one single SOAP dimension accounting only of the total dataset’s variance (neglected and discarded in typical variance-based dimensionality-reduction approaches) allows resolving a remarkable amount of information, classifying/discriminating the bulk of water and ice phases, as well as two solid-interface and liquid-interface layers as four statistically distinct dynamical molecular environments. Adding more dimensions to this one is found not only ineffective but even detrimental to the analysis due to recurrent negligible-information/non-negligible-noise additions and “frustrated information” phenomena leading to information loss. Such effects are proven general and are observed also in completely different systems and descriptors’ combinations. This shows how high-dimensional analyses are not necessarily better than low-dimensional ones to elucidate the internal complexity of physical/chemical systems, especially when these are characterized by non-negligible internal noise.
@@ Introduction
A common assumption when studying complex systems (of any scale), about which little is known a priori, is that high-dimensional analyses are desirable, and even necessary, to avoid losing information due to prior assumptions or incomplete characterization. However, selecting a set of physically relevant descriptors capable of providing a complete characterization of a specific system requires a deep knowledge of the system itself, which is not always available. Moreover, managing high-dimensional analyses and extracting meaningful physical insights from them is often challenging. Fundamental questions arise: What is the intrinsic dimension of the dataset [facco2017estimating, dinoia2024beyond]? Which components are the most relevant?
@@
Similarly to other descriptors, such as atomic cluster expansions (ACE) [drautz2019atomic] or N-body iterative contraction of equivariants (NICE) [nigam_recursive_2020], SOAP provides high-dimensional mathematical representations of local density and order/disorder. Recently, SOAP has been successfully applied to investigate complex molecular systems, including liquid-solid and liquid-liquid coexistence in aqueous systems [Donkor2024spatavg, offeidanso2022, Ansari2020, Martino2024], and non-trivial dynamics and local disorder events in metal surfaces and nanoparticles [Caruso2023tsoap, Caruso2024, crippa2023machine, Rapetti2023, Cioni2024, perrone2024], and ionic behaviors in dense environments [Lionello2022].
@@
This finding raises several important questions: To what extent is high-dimensionality truly necessary to capture the physics of these systems? How can a high-dimensional analysis be less informative than its dimensionally-reduced counterparts? What type of information remains hidden, where is it located, and how can it be extracted? Is the information contained in the orthogonal dimensions of a high-dimensional feature space always additive, does adding a second dimension always contribute new information? These questions are particularly central when using intrinsically high-dimensional descriptors such as, e.g., SOAP or ACE [drautz2019atomic], but they also apply to any combination of descriptors, whether physics-inspired, abstract, or general. At their core, these questions challenge a fundamental issue in physics: How to effectively extract meaningful information from high-dimensional analyses, a problem that is central in many fields.
@@
Most often, high-dimensional dataset extracted from, e.g., molecular dynamics (MD) trajectories (but this is true for any type of trajectory, including experimentally-resolved ones), are studied under an ergodic assumption as a collection of temporally-independent data, and clustering methods are used to detect high-density peaks identifying dominant/recurrent SOAP domains [Capelli2022, Gardin2022, Gasparotto2019]. However, we have recently shown that significant information may be hidden in the temporal correlations of the data – such as local, sparse but dominant fluctuations, important local transitions, etc. – that are typically overlooked in pattern recognition approaches which neglect time correlations. We have also developed an efficient method capable of maximizing the extraction and classification of information in time-series data, systematically distinguishing statistically relevant fluctuations from noise as a function of the time-resolution used in the analysis [Becchi2024layer].
@@
In this work, we leverage such approaches to tackle fundamental questions relevant in high-dimensional analyses of molecular (and non-molecular) systems, such as: To what extent a high-dimensional analysis approach is really beneficial compared to lower-dimensional ones? How to compare information quantity vs. information quality (relevance)? As a first prototypical example of an information-rich dataset, we analyze SOAP data extracted from atomistic MD simulation of a periodic box where liquid water and ice coexist in dynamical equilibrium at the melting temperature [Caruso2023tsoap, crippa2023machine, Becchi2024layer, Caruso2024, Martino2024]. This provides us with a prototypical system in equilibrium that respects a detailed balance and, at the same time, exhibits non-trivial internal complexity. This is explored here by computing 576-dimensional SOAP spectra for each of the 2048 water molecules in the system every ps (sampling interval) over 50 ns of MD trajectory. This results in a dataset of 576-dimensional SOAP spectra that we analyze both by ignoring or considering the time correlations between the data (i.e., pattern recognition vs. time-series analyses).
@@
Our results demonstrate how dimensionality reduction approaches, such as Principal Component Analysis (PCA), fall short when dealing with noisy data. Similar approaches allow discriminating and classifying only two solid ice and liquid water environments: i.e., the two most populated dominant phases in the system (recently Time-lagged Independent Component Analysis(tICA) [molgedey1994, perez2013] analyses were also proven not much more instructive).[Caruso2024] On the other hand, we show how time-series data contained in a single SOAP component, which accounts remarkably only of the total variance of the SOAP dataset (and that is thus typically discarded in PCA), can provide more information than that attainable by analyzing all dimensions in the SOAP dataset altogether. Our results reveal that adding more (noisy) components to this single one can be not only non-beneficial but also detrimental to the analysis, a phenomenon we describe as “noise-frustrated information”. Finally, we analyze a completely different system – whose trajectories are obtained experimentally and are analyzed by combining physics-based descriptors – showing that such effects are not exclusive of SOAP data, but can arise in any noisy multidimensional analysis. The findings presented here challenge classical paradigms in data analysis, highlighting the critical importance of focusing on information quality (relevance) rather than quantity to minimize information loss and misinterpretation.
@@ Results
The extraction of information from the trajectories of such a simulated system involves three main steps: (i) acquisition of raw data, which consist in the sequence of the particles’ coordinates acquired every ps, over ns of MD simulation; (ii) selection of an appropriate descriptor to convert the trajectory of each molecule – given by the time-series data – into a dataset ; (iii) analysis of the dataset to extract relevant information on the system’s physics (see Figure 1b). The information that can be effectively collected from the system depends on the choice of the descriptor in step (ii). The correct choice of descriptor is crucial for two main reasons. First, different descriptors capture distinct aspects of the system’s behavior: for example, a structural descriptor would be useless to study purely dynamical features [crippa2023machine]. Second, each descriptor has an intrinsic signal-to-noise ratio, which influences the clarity and reliability of the extracted information [Martino2024]. When the physics of a system is unknown, selecting a general, agnostic – possibly intrinsically high-dimensional – descriptor would be desirable in order to minimize the risk of introducing assumptions that could bias the analysis or result in incomplete information.
@@
It is particularly noteworthy that, in Figure 2a, four PCs were required to reach of the cumulative variance, implying that the dataset was approximately four-dimensional. However, the local noise reduction allows PC1 alone to account for nearly the entire variance of the dataset, indicating that the “relevant” information, without noise, is essentially mono-dimensional. This suggests that the high-dimensionality observed in the original dataset is primarily a reflection of noise, not of the inherent complexity of the system.
@@
Information quantity vs. information quality in noisy datasets
@@
Notably, all our evidences demonstrated that this phenomenon is general and can occur independently on the number of dimensions, their individual noise levels (native vs. denoised), or the dimensionality reduction approach employed in the analysis. To further illustrate this point, Supporting Figure SI7 shows the manifold representations of the SOAP and SOAP datasets (computed using UMAP [mcinnes2018umap, Donkor2024spatavg]). This approach operates directly on the full high-dimensional dataset – meaning combining all component information “correctly” – and captures comparable physically relevant information to the one detected by the single component (Figure 4f, g). This confirms that discriminating the components by their variance is not necessarily the best approach, and better metrics should be used to extract the relevant information [Ebrahimi1999]. Similar to PCA, Time-lagged Independent Component Analysis (tICA) [molgedey1994, perez2013] performed on the same system, combined with KMeans clustering, also allowed to discriminate ice from water, struggling to resolve other dynamical microscopic domains (such as the ice-water interface).
Information frustration effects are also observed when additional data are added in time (by increasing the sampling frequencies) rather than in space (by adding SOAP components). To investigate the effects of different sampling frequencies, we compared the results obtained analyzing the MD trajectories (raw data) sampled every ps (Figures 2,3,4) with those obtained by increasing or decreasing the sampling frequency. We therefore resampled the 50 ns MD trajectory at intervals of 4 ps, 20 ps, 80 ps, and 100 ps. The expectation is that Onion Clustering, which extracts relevant information about physical events in the system (for instance, the detectability of the ice/water interface environment up to a certain ), should maintain or enhance its ability to resolve this environment when using finer sampling ( ps). Conversely, coarser sampling ( ps) might compromise the resolution, potentially preventing proper detection of the interface. However, the results of these analyses reveal issues analogous to those encountered when adding more spatial information. As shown in Figure 5b, the black line represents the reference case ( ps): the third environment – the ice/water interface – is detectable within the resolution range 2 ns ns in the PC1 time-series, as seen in Figure 2c. Increasing the sampling time to 4 ps or 20 ps produces striking effects, as the temporal window within which the interface is detected shifts backward in the time-series. For a 20 ps sampling interval (which doubles the number of frames compared to the reference), the detectable temporal window for the interface shifts to 1 ns ns (green data). With a 4 ps sampling interval (a tenfold increase in the number of analyzed frames), the interface is detectable only between 200 ps and 300 ps, a much shorter period compared to the nanosecond-scale range observed in other cases.
@@
Generality of the concept: Quincke rollers as a second different case study
@@
In the tests of Figures 1-5 we used SOAP as a representative example of a high-dimensional descriptor, providing high-dimensional datasets. However, the same phenomena can occur independently of the specific system or the set of descriptors/variables selected. To illustrate the generality of these concepts, we analyzed a completely different system, characterized by unique complex internal dynamics. In this case, we examined time-series obtained from experimentally-acquired trajectories of a system involving the two-dimensional collective motions of colloidal polymeric microparticles, known as Quincke rollers. Readers seeking more details on this specific system are referred to Refs. [Liu2021quinckerollers, Becchi2024layer, Caruso2024]. Briefly, Quincke rollers are dielectric polystyrene particles (m in diameter) suspended in a weakly conductive fluid and confined within a two-dimensional cell. Under the influence of a weak vertical DC electric field, these particles can exhibit complex collective motions, such as vortexes and waves. While this is a higher-scale system, completely different from the ice-water molecular-scale one studied in the previous sections, it is useful to show how also in this case we encounter similar ”information frustration” effects/issues, which are thus not limited to certain scales, systems, or descriptors, but are general in high-dimensional analyses of noisy systems.
Here, we analyze the trajectory we resolved for one example Quincke rollers system [Becchi2024layer], consisting of 6921 particles moving within a m cell over a real-time duration of 0.25 s (310 frames in total). During this interval, a collective density wave is seen to propagate from left to right in the microscopy field (see Figure 6a-b).
@@
For the sake of generality, in this case our analysis uses two human-based descriptors. Recently, we have demonstrated that integrating static descriptors, such as SOAP, with dynamic ones, such as LENS, can enrich high-dimensional analyses [crippa2023detecting]. In the same spirit – and reflecting common practice in studying collective dynamical events in active matter [peruani2008mean] – we selected two descriptors: (i) the minimum neighbor distance , as a proxy for the local particle density, and (ii) the local alignment of the particle velocity . This choice results in a bi-dimensional analysis (details are provided in the Methods section). We analyzed the denoised mono- and bi-dimensional time-series of the 6921 particles. Using Onion Clustering, was identified as a less informative descriptor (Figure 6c). From its time-series, two environments are detected up to ms: one representing the wave’s core (having higher density) and the other identifying static particles (with lower density). In contrast, the particles’ local velocity alignment (Figure 6d) is found to be more informative, capturing the system’s collective motion and identifying three distinct environments for resolutions higher than ms, and two environments for ms. When three clusters are detected, Onion Clustering distinguishes the wave’s core (Figure 6d in fuchsia) and two regions (cyan) where particles transition between static and dynamic states around the wave’s core. These regions are shown in the snapshot in Figure 6d, alongside with the static particles’ environment (dark violet). For ms, Onion Clustering fails to resolve any statistically distinct dynamical cluster, indicating that, in the system, all particles change environment at list once during the trajectory. The residence time of particles in any cluster is therefore less than 140 ms.
@@ Conclusions
As a prototypical case study, we analyzed Smooth Overlap of Atomic Position (SOAP) spectra [bartok2013representing] extracted from MD simulations of coexisting ice and liquid water at dynamic equilibrium at the solid/liquid transition temperature. Typically, SOAP analyses of such systems involve dimensionality reduction methods, such as Principal Component Analysis (PCA) followed by clustering based on a combination of principal components (PCs) that account for high cumulative variance. However, approaches like this ignore the time-correlation in the data; this can result in significant information loss and limit the ability to capture key dynamic events [Becchi2024layer]. To maximize the information extraction, we opted to analyze the SOAP data in time-series using Onion Clustering [Becchi2024layer], which allows resolving the information within individual SOAP PC1 and PC2 time-series, as well as their combination in bi-variate time-series. Our results demonstrate that two-dimensional analysis systematically lead to information loss, independently of whether the time-series data are noisy or denoised [Donkor2024spatavg]. This phenomenon persists even when analyzing two of the original SOAP components, such as 63 and 237 (see Figures 4-5), and thus regardless of the chosen dimensionality reduction approach. We call this effect “frustrated information”, wherein relevant information captured in one dimension manifest as noise in another one.